OpenStudy (anonymous):

Please help me with this question....Convert 3x + 2y = 5 into slope-intercept form..

OpenStudy (anonymous):

subtract 3x from both sides then divide by 2

OpenStudy (anonymous):

slope intercept form is y=mx+b. So, isolate y. 2y=-3x+5. Divide 2 on both sides. y=-3/2x+5/2

OpenStudy (anonymous):

Thank you